OpenStudy (anonymous):

Choose the equation of the line passing through the point (4, -2) and perpendicular to y = one thirdx - 7.

OpenStudy (anonymous):

the slope formula for the line perpendicular to the line y= 1/3 x -7 is m2=-1/m1 then what is the slope of the line perpendicular to it? is it m=1/3 or m=-3? then use the point (4, -2)=(x1,y1), and the formula y-y1=m(x-x1) to find the equation of the line perpendicular to it

OpenStudy (anonymous):

@ 0214